During report, the nurse who is leaving tells the oncoming nurses that she just hung a liter of IV fluid for a patient...

  1. During report, the nurse who is leaving tells the oncoming nurses that she just hung a liter of IV fluid for a patient with mild heart failure, and that is to be run over 12 hours. When the nurse enters the room 30 minute later, he finds 600 mL left in the bag. Because he is concerned about the patient too much fluid, he calls the physician, who tells him to slow down the IV so that the remaining volume infuses over the next 11 hours. What rate will the nurse need to set the pump at to accomplish this?   (Answer in mL per hour)  

  1. To prepare for a colonoscopy, the patient will need to drink 1.8 Liters of a prep solution. The patient is instructed to drink 400 mL initially, then the rest over the next 6 hours.

How many mL per hour will the patient need to drink to consume all of the prep solution?  

The patient asks the nurse how much this is in ounces or cups. How much is it in each?
Answer #1

Ans 1:

Given: Volume =600 ml

Time=11 hours

To find: Flow Rate =?

Formula:

Flow rate for infusion pump=Volume in ml /Time in hours

Calculation: 600 ml /11 hrs=54.545 ml/hr

The nurse should set the infusion pump at 55 ml/hr rate to accomplish this.

Ans: 2

total solution is 1.8 litre which is 1800 ml.( as 1 litre equals to 1000 ml therefore 1.8 litre will be 1800 ml)

patient initially already consumes 400 ml therefore 1400 ml solution is remained in bottle.

this 1400 ml to be consumed in 6 hrs

therefore desireed requireed solution=volume/time

1400/6=233.33

therefore patient should drink 233 ml per hour to consume all of the prepared solution.

1 ounce is equal to 30 ml approximately (round figure).

therefore initially 13.33 ounces followed by 7.8 ounces to be drank every hourly.( i-400/30=13.5 ounces , 233.33/30=7.77)